I would like to know how do I add, subtract, multiply, and divide NEGATIVES AND POSITIVES FRACTIONS. AND WHAT IS A NEGATIVE + NEGATIVE=
, NEGATIVE + POSITIVE, POSITIVE +NEGATIVE, NEGATIVE x NEGATIVE, POSITIVE x NEGATIVE. NEGATIVE DIVIDE POSITIVE

Answer by elima(1433)

(+)+(+)=(+)
(+)+(-)=subtract and take the sign fro the LARGER #
ex;
5+(-3)=2
2+(-6)=-4
(-)+(-)= (-)
(-)-(-)=change the minus sign to a negative, and the rules for adding (+)and(-) apply;
ex;
-4-(-5)=-4+5=1
-5-(-3)=-5+3=-2
Multiplying and dividing negatives and positives are much easier than adding and subtracting;
(+)*(+)=(+)
(-)*(-)=(+)
(+)*(-)=(-)
Same rules apply for dividsion
Hope this helps
=)